OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -10 and 2x <(with line under) 12

OpenStudy (anonymous):

Do you have a screenshot I can look at so I can take a look at your answer choices?

OpenStudy (anonymous):

\[x - 2 > -10 \quad and \quad 2x \le 12\] First solve for x in both equations. \[x - 2 > -10\]Add 2 to both sides.\[x > -8\] \[2x \le 12\]Divide both sides by 2.\[x \le 6\]

OpenStudy (anonymous):

A number line with an open circle on -6, shading to the left, and a closed circle on 9, shading to the right. A number line with a closed circle on -6, shading to the left, and an open circle on 9, shading to the right. A number line with an open circle on -6, a closed circle on 9, and shading in between. A number line with a closed circle on -6, an open circle on 9, and shading in between.

OpenStudy (anonymous):

@RobZBHayes

OpenStudy (anonymous):

So to graph remember: Less/Greater than(< or >)... Open circle. Less/Greater than EQUAL to (<= or >=)... Closed circle Less than (< or <=)... Shading to the left. Greater than (> or >=)... Shading to the Right.

OpenStudy (anonymous):

Yea i know that but idk how to solve the equation

OpenStudy (anonymous):



OpenStudy (anonymous):